OpenStudy (anonymous):

can someone help me with finding the equation of the tangent line to the graph of the function at the indicated point? y=x^3*squareroot(x^3+3) at x=1

OpenStudy (anonymous):

I know I can rewrite it first as y=x^3(x^3+3)^1/2

OpenStudy (anonymous):

take the derivative replace \(x\) by 1, i.e. find \(f'(1)\) that is your slope then use the point slope formula

OpenStudy (anonymous):

if i take the derivative would i get y=(3x^2)*1/2(X^3+3)^-1/2

OpenStudy (anonymous):

you have \[x^3\sqrt{x^3+3}\] derivative requires product and chain rule get \[f'(x)=3x^2\sqrt{x^2+3}+x^3\frac{3x^2}{2\sqrt{x^3+3}}\]

OpenStudy (anonymous):

clean up with some algebra if you need, or just replace \(x\) by \(1\) to get the slope

OpenStudy (anonymous):

oh okay i tried doing it with just the product rule before but i never applied the chain rule okay

OpenStudy (anonymous):

the slope would be ... let me work it out sorry i am little slow

OpenStudy (anonymous):

slope is 21/4?

OpenStudy (amistre64):

\[f'(x)=3x^2\sqrt{x^2+3}+x^3\frac{3x^2}{2\sqrt{x^3+3}}\] \[f'(1)=3\sqrt{4}+\frac{3}{2\sqrt{4}}\] \[f'(1)=6+\frac{3}{4}\]

OpenStudy (anonymous):

im working it out again i think i missed a number in missed a number

OpenStudy (anonymous):

yeah i added a one somewhere before

OpenStudy (amistre64):

calculus is easy, its the algebra that tends to throw ya ;)

OpenStudy (anonymous):

yeah when i get stuck and then ask for help it usually doing the algebra to fast that gets me

OpenStudy (anonymous):

so to find the equation of the line can i plug in the x value of 1 and slope of 27/4 and solve for b? oh but i need to plug in x to the original equation first right? to get he y value?

OpenStudy (amistre64):

there are many "forms" of a line equation. The 2 most common forms I see associated with a tangent line are:\[y-f(a)=f'(a)(x-a)\]or\[y=f'(a)(x-a)+f(a)\] given that x=a=1 in this case, and f'(1)=27/4 ... yes, you will need to determine y=f(1) to finish it up

OpenStudy (amistre64):

it might be useful to note that:\[y=f'(a)(x-a)+f(a)\\y=f'(a)x-f'(a)a+f(a)\] \[y=f'(1)x-f'(1)+f(a)\]would be the usual y=mx+b from algebra

OpenStudy (amistre64):

f(1), not f(a) ... lol, thats what i get for trying to look smart

OpenStudy (anonymous):

it's okay your smart i'm the one that is not so smart today. but...

OpenStudy (anonymous):

i though i would plug in the slope to the original equation to determine a value for y and then plug in x, y and the slop to get be and then write my new equation. but i'm not getting an answer on my sheet that's an option

OpenStudy (amistre64):

f(x) = y = x^3*sqrt(x^3+3); x=1 f(1) = y = 1^3*sqrt(1^3+3) = sqrt(4) = 2 so the given point is (1,2) for the line equation to work out ;)

OpenStudy (anonymous):

yes and then i got -19/4 as b

OpenStudy (anonymous):

i made an algebra mistake again. that's why i usually do everything twice.

OpenStudy (amistre64):

\[y-2=\frac{27}{4}(x-1)\] \[y=\frac{27}{4}x-\frac{27}{4}+2\] \[y=\frac{27}{4}x-\frac{27+8}{4}\]

OpenStudy (amistre64):

the first one is a usual tangent line format .. point slope form the last one is it swapped about into slope intercept form

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

youre welcome, ive got real world job stuff to take care of for a little bit ... so i wont be around :) but there are plenty of people here that are smarter than me to help you out good luck

OpenStudy (anonymous):

well i really appriciate it and have a good day